# Dynamics of Clusters of Anyons in Fractional Quantum Hall Fluids

### v1

@@ IV.2 LDOS Measurement for Non-Abelian Moore-Read Phase
In the MR phases, the $\Delta M=0$ state is even more deeply buried in the continuum of gapped excitations, regardless of the e-e interaction being long-range 1LL Coulomb or comparatively short-range 1LL Yukawa potentials. Across all simulated LDOS results for both odd and even $N_{e}$ systems under various e-e interactions, the most experimentally detectable lowest peaks under the assumption of a small bias voltage correspond to states with very large relative angular momentum within the MR anyon cluster. Consequently, the tunneled states observed in experiments are likely those with low energies and relatively larger $\Delta M$ (i.e. larger anyon-cluster sizes), rather than the $\Delta M=0$ state. However, depending on the strength of the bias voltage and the details of the STM tip, it does not necessarily correspond to the absolute lowest energy state.
Notably, even for MR phases where the energy spectra appear similar under various realistic electron-electron interactions as we mentioned in the last section, the LDOS remains sensitive to these variations, manifested from the strengths and locations of the density peaks in Figure 6. This underscores STM’s potential as a tool for detecting and verifying effective anyon interactions in both Abelian and non-Abelian FQH systems. However, compared to the Laughlin state—where peaks are well separated even at lower resolutions—MR systems require measurements of higher resolution. This is understandable since removing one electron creates three Laughlin quasiholes, but four MR quasiholes, and the latter leads to many more configurations with similar energies.
@@ V Conclusion
In this paper, we investigated the effective interactions of anyons in Laughlin and Moore-Read FQH phases derived from diverse electron-electron interactions, both model and realistic ones. The effective interaction within anyons is an interpolation between fermions’ repulsive and bosons’ attractive force, exhibiting various and unusual behaviors depending on the detailed forms of the electron-electron interaction in the 2DEG. The recently developed STM experiments in FQH systems could be a new platform to explore these properties of anyons. Previously only the most tightly bound anyon clusters are considered in such experiments, which we show have very high energy and many other bound anyon states will be involved in tunneling at lower energies. We point out that effective anyon interactions are important factors in understanding the LDOS results, which were ignored by several previous theoretical studies. Since anyons can be considered the elementary particles for the corresponding FQH phases and are potentially useful for the storage and processing of quantum information, the interplay between their effective interactions and the correlated topological phases should be considered as a significant feature when investigating their properties both theoretically and for future applications.
@@ Acknowledgements.
Some of the numerics in this work are performed using DiagHam package for which we would like to thank all the authors. This work is supported by the National Research Foundation, Singapore, under the NRF Fellowship Award (NRF-NRFF12-2020-005), Singapore Ministry of Education (MOE) Academic Research Fund Tier 3 Grant (No. MOE-MOET32023-0003) “Quantum Geometric Advantage”, and Singapore Ministry of Education (MOE) Academic Research Fund Tier 2 Grant (No. MOE-T2EP50124-0017).
Supplementary Materials
.1 Energy spectrum for three-anyon clusters in Laughlin 1/3
Figure S1 (a) and (b) show the effective interacting energy, whose definition has been explained in the main text, within a three-anyon cluster in Laughlin 1/3 FQH phase introduced by model Hamiltonian $\hat{V}_{int}=t\hat{V}_{1}+\hat{V}_{2n+1},\ n=1,2$ ($\hat{V}_{3}$ and $\hat{V}_{5}$). $t\gg 1$ is required to make sure that we are inside the $\mathcal{H_{L}}$. Here the number of quasiholes $N_{qh}=3$. For a multi anyon system ($N_{qh}\geq 3$), there are two-body interactions in-between each pair of anyons while the CHS also allows the existence of effective many-body interactions. So the constituents of the effective interaction in a three-anyon cluster are three pieces of two-body interactions and the three-body interaction among the whole cluster. We will discuss the entire effective interaction as a whole and not dig into details to separate and clarify the effects from its different components in this work.
In both cases for $\hat{V}_{3}$ and $\hat{V}_{5}$, the lowest energy state is $\Delta M=6$ state where the effective interactions are negative, representing an energetically favored $\Delta M=6$ three-anyon cluster. The finite-size scaling of the value for this effective interacting energy at $\Delta M=6$ in Figure S1 (c) verifies this negativity at thermal dynamic limit.
Figure S1: Effective interaction of 3-anyon cluster under model Hamiltonian (a) $\hat{V}_{3}$ and (b) $\hat{V}_{5}$ in $\mathcal{H_{L}}$. (c) Finite-size scaling of the interaction energy with model Hamiltonian $\hat{V}_{3}$ (blue) and $\hat{V}_{5}$ (purple) at $\Delta M=6$ in a cluster of three Laughlin 1/3 anyons.
.2 Self Energy of MR $e/4$ quasiholes
Figure S2: Finite-size scaling of the self energies of the MR quasiholes induced by PPs.
Inserting a single flux quantum into the MR ground state creates two $e/4$ MR quasiholes, a consequence of flux fractionalization in non-Abelian phases. The total energy of the system then have three contributions: the ground state energy $E_{g}$, twice the self-energy of a single $e/4$ MR quasihole, and the interaction energy between the two anyons. Unlike the Laughlin case, the self-energy cannot be extracted directly due to the coexistence of interactions between the anyons. However, by placing the two quasiholes at the North and South poles of the sphere, their interaction energy decays to zero as the system size increases and the separation becomes sufficiently large.
The effective energy associated with the quasiholes and their interaction is defined as
$E_{\text{eff}}=N_{qh}E_{\text{qh}}+E_{\text{int}}=E_{\text{total}}-E_{\text{g}},\quad N_{qh}=2$ (S1)
where $E_{\text{total}}$ is the total energy of the system, $E_{qh}$ is the self-energy of a single $e/4$ quasihole, and $E_{\text{int}}$ denotes the interaction energy between the quasiholes. The self-energy of a single quasihole is then approximated by extrapolating the effective energy in the thermodynamic limit:
$E_{\text{qh}}=\lim_{N_{e}\to\infty}\frac{1}{2}E_{\text{eff}}$ (S2)
It is assumed that the self-energies of the two anyons in a system with a fixed topological number (“$1$” or “$\psi$”) are identical. However, since the self-energy may differ between even and odd $N_{e}$ systems, the two cases are treated separately.
Figure S2 shows $E_{\text{eff}}/2$ plotted as a function of $1/N_{e}$ for odd and even $N_{e}$ systems with various model electron-electron interactions. The data points are extrapolated to extract the self-energies according to Eq.(S2), using the largest system sizes accessible within our numerical limits–$N_{e}=19$ for odd-parity and $N_{e}=20$ for even-parity systems. We note that the extraction of self-energies is strongly constrained by the system-size limitations of exact diagonalization. Since larger systems yield weaker interactions and thus more reliable data, the fitting and extrapolation are weighted more heavily toward results from larger systems.
Additionally, the behaviours of $E_{\text{eff}}$, and consequently the self-energies of the $e/4$ MR quasiholes, exhibit a more stable and systematic trend in odd-$N_{e}$ systems (topological quantum number “$\psi$”) compared to even-$N_{e}$ systems (topological quantum number “$1$”). This also implies that the neutral Majorana fermion mode exhibits a more stable self-energy than the charged U(1) bosonic mode according to the bulk-dedge correspondance. The electrically neutral Majorana mode is less sensitive to interactions and finite-size effects Read and Moore (1991); Read and Green (2000), whereas the charged U(1) sector couples strongly to background potentials and orbital discretization Rezayi and Haldane (2000). Moreover, the Majorana mode is topologically protected as part of the non-Abelian order, reducing its susceptibility to perturbations Read and Moore (1991); Wen (2004). Similar conclusions regarding the robustness of neutral edge modes have been discussed in studies of edge structures Milovanović and Read (1996).
.3 Energy spectrum for four-anyon clusters in Moore-Read phase
Figure S3 shows the energy spectrum of quasihole excitations for systems with odd ($N_{e}=13$, top row) and even ($N_{e}=14$, bottom row) parities with electron-electron model interaction $\hat{V}_{1}$, $\hat{V}_{3}$, and $\hat{V}_{5}$ inside the MR CHS. Across all cases, the lowest-energy states are located at relatively large $\Delta M$. The counting, which is one of the reflection of the topological property of the system, is distinct in between systems hold the two different paritiesYang (2021).
As a transition from model Hamiltonian to realistic ones, we studied the low-lying energy spectrum of MR four-quasihole systems with the model Hamiltonian Eq.(S3) in 1LL.
$\hat{H}_{int}=t\hat{V}_{3}^{3bdy}+\hat{V}_{\text{Coulomb}}$ (S3)
where $\hat{V}_{3}^{3bdy}$ is the three-body pseudopotential that defines the MR conformal Hilbert space $\mathcal{H_{MR}}$. When $t\to\infty$, the system is entirely within the $\mathcal{H_{MR}}$. The Hilbert space will gradually evolve to the real space as $t$ is tuned down.
Figure S4 shows the evolution of the four-quasihole excitation spectra as $t$ decreases, for both odd (top row) and even (bottom row) $N_{e}$ systems. Remarkably, the structure of the quasihole spectrum remains largely unchanged even at small $t$, despite the dressing from neutral excitations outside the $\mathcal{H_{MR}}$, as indicated by the little variation in the distribution of the quasihole excitation states and their dependence on the relative angular momentum $\Delta M$.
The energy gap $\Delta E$ between the quasihole excitation manifold and the continuum as a function of $t$ is shown in Figure S5. For both $N_{e}=13$ and $N_{e}=14$ systems, the gap exhibits a almost perfectly linear dependence on $t$ and remains open until $t$ is reduced to a very small value. This indicates that the quasihole structure within the MR conformal Hilbert space is largely preserved across a wide range of $t$. And only when $t$ becomes sufficiently small, where the system effectively approaches the full realistic Hilbert space limit, does the topological gap finally collapse, signaling the breakdown of the gapped phase.
Figure S3: Energy spectrum of 4-anyon cluster under model Hamiltonian $\hat{V}_{1}$, $\hat{V}_{3}$, and $\hat{V}_{5}$ in both odd and even $N_{e}$ systems
Figure S4: Energy spectrum of the model Hamiltonian. Red crosses denote the quasihole excitation states while the blue symbols correspond to higher-energy states outside the $\mathcal{H_{MR}}$. First row: odd, second row: even. t decreases from left to right: (a) t =1, (a) t =0.5, (a) t =0.1, (a) t =0.05.
Figure S5: Energy gap between quasihole excitations and other parts in the spectrum of the four-quasihole MR phase with model Hamiltonian S3 plotted against t.
.4 Spatial extent analysis
The estimated radius of a few-anyon cluster $R_{1}$ (in units of $l_{B}$) from the first moment is given by:
[…]
where $\delta\rho(\theta)=\rho(\theta)-\rho_{0}$, and $\rho_{0}$ is a background density. We take $\rho_{0}$ to be the ground-state density without additional fluxes in our analysis. The second moment $R_{2}$ in units of $l_{B}$ is:
[…]
Eq. (S4) and (S5) give two model-independent estimate of the arc lengths for a few-anyon cluster on sphere. The arc length will asymptotically approach the radius of the anyon cluster on planar geometry in the thermodynamic limit.
.4.1 Laughlin
The two radii for the $\Delta M=12$ three-anyon cluster with electron-electron Coulomb interaction, computed in $N_{e}=11$ system, are:
$R_{1}^{l}\approx 6.28l_{B},\ R_{2}^{l}\approx 6.92l_{B}$ (S6)
Typically, the magnetic length is estimated as $l_{B}\sim 26\text{nm}/\sqrt{B}$, where $B$ is the strength of the magnetic field in Tesla. The Laughlin 1/3 state is realized at magnetic fields of $B\sim 10-15\text{T}$ in experiments. These values provide an estimate for the spatial extent of the anyon cluster as: $R_{1}^{l}\sim 40-50\text{nm}$, $R_{2}^{l}\sim 45-60\text{nm}$. Similarly, for the Yukawa interaction with $\lambda=1$, the lowest-energy $\Delta M=6$ state has the first and second moments:
$R^{l^{\prime}}_{1}\approx 6.38l_{B},\ R^{l^{\prime}}_{2}\approx 7.23l_{B}$ (S7)
which correspond to the spatial extents: $R^{l^{\prime}}_{1}\sim 40-50\text{nm}$, and $R^{l^{\prime}}_{2}\sim 45-60\text{nm}$.
The calculated spatial extents of anyon clusters from the first and second moments are unexpectedly similar under both Coulomb and Yukawa interactions. This similarity arises because the moment-based calculation primarily captures the weighted average separation of the density, where both the core region and oscillatory tails contribute significantly. Although the density profile under Yukawa interaction exhibits core dominance, while the Coulomb profile features a more oscillatory structure, the global contribution from the entire density profile ultimately yields comparable spatial extents (for density profiles and more detailed discussions, see the Supplementary). Thus, the spatial extent calculation reflects the global density distribution, while $\Delta M$ more accurately captures the cluster compactness and inside structures.
The density profile of three-anyon clusters under electron-electron Coulomb interaction in Laughlin phase exhibits multiple sharp peaks, indicating long-range correlations and oscillatory structures (Fig S6 (a)). This pattern reflects weak central binding, with the cumulative contribution rising gradually and reaching the 50% mark at a large polar angle ($\theta\approx 2.0$ radians), suggesting that the density deviation is more evenly spread over the sphere without a dominant core. In contrast, the short-range Yukawa interaction produces a density profile with a prominent central peak and mild, smooth oscillations extending toward the edges (Fig S6 (b)). This indicates that the central region holds a significant fraction of the density deviation, reflecting strong short-range binding. The cumulative contribution shows that the core region (within $\theta\approx 1.5$ radians) already accounts for about 50% of the total density deviation. While the core is compact, the long, smoothly decaying tail contributes notably, leading to a larger-than-expected radius.
The moment-based calculation reflects the global density distribution, giving substantial weight to both the central peak and tail contributions. This approach does not fully capture the difference in cluster compactness or the detailed structures arising from different electron-electron interactions. Although the Yukawa interaction forms a tightly bound core (as indicated by both the density profile and the reduced relative angular momentum $\Delta M$), the long tails result in a spatial extent comparable to that of the Coulomb case.
In summary, based on the finite-system study we perform here, the Coulomb density profile is more oscillatory and spread out, while the Yukawa profile shows core dominance with smooth tails. The cumulative contribution plots further demonstrate the differences in spatial distribution, emphasizing the need to distinguish between compactness and global spatial extent when analyzing anyon cluster sizes. Although the calculated spatial radii may appear similar, the underlying clustering mechanisms differ significantly with the two different electron-electron interactions. The spatial extent calculation primarily reflects the averaged density distribution, while $\Delta M$ will capture the internal correlation and cluster compactness.
Figure S6: Quasihole density for (a) the $\Delta M=12$ cluster with Coulomb interaction, and (b) Yukawa interaction ($\lambda=1$). $N_{e}=12,N_{o}=36$.
.4.2 Moore-Read
The quasihole density of MR system containing a four-quasihole cluster in odd and even $N_{e}$ ($N_{e}=15,16$ respectively) systems are shown in Figure S7. They exhibit similar oscillatory structures with multiple sharp peaks, indicating the long-range structure of the four-anyon cluster with 1LL Coulomb interaction. The radii of the loose anyon cluster are estimated by Eq. (S4) and (S5), similarly as in Laughlin’s case. For the lowest energy state as $\Delta M=12$ in $N_{e}=15$ (odd) system, the two radii are:
$R_{1}^{MR}\approx 6.77l_{B},\ R_{2}^{MR}\approx 7.40l_{B}$ (S8)
And for the $\Delta M=14$ anyon cluster in $N_{e}=16$ (even) system,
$R_{1}^{MR^{\prime}}\approx 5.23l_{B},\ R_{2}^{MR^{\prime}}\approx 5.92l_{B}$ (S9)
Experimentally, the $\nu=5/2$ state (partial filling $\nu^{*}=1/2$ in the 1LL) is realized when the magnetic field $B\sim 2-3T$. These radii correspond to the spatial extents: $R_{1}^{MR}\sim 100-125\text{nm}$, $R_{2}^{MR}\sim 110-140\text{nm}$; $R_{1}^{MR^{\prime}}\sim 75-95\text{nm}$, $R_{2}^{MR^{\prime}}\sim 85-110\text{nm}$. The spatial extent of the anyon clusters with the lowest energy in the MR phase exhibits a difference between odd and even electron systems, suggesting a parity effect in the cluster configuration in the finite systems. And the predicted spatial extents are much larger compared to the Laughlin’s cases.
Figure S7: Quasihole density under 1LL Coulomb interaction for (a) $N_{e}=15$ and (b) $N_{e}=16$ systems, both containing four MR quasiholes.

### v2

@@ IV.2 LDOS Measurement for Non-Abelian Moore-Read Phase
In the MR phases, the $\Delta M=0$ state is even more deeply buried in the continuum of gapped excitations, regardless of the e-e interaction being long-range 1LL Coulomb or comparatively short-range 1LL Yukawa potentials shown in Figure 4. Across all simulated LDOS results for both odd and even $N_{e}$ systems under various e-e interactions, the most experimentally detectable lowest peaks under the previous assumption correspond to states having very large relative angular momentum within the MR anyon cluster. However, depending on the strength of the bias voltage and the details of the STM tip, it does not necessarily correspond to the lowest energy state.
Notably, even for MR phases where the energy spectra appear similar under various realistic electron-electron interactions, as we mentioned in the last section, the LDOS remains sensitive to these variations, manifested from the strengths and locations of the density peaks in Figure 6. This underscores STM’s potential as a tool for detecting and verifying effective anyon interactions in both Abelian and non-Abelian FQH systems. However, compared to the Laughlin state—where peaks are well separated even at lower resolutions—MR systems require measurements of higher resolution (for example, FWHM = $0.001\ e^{2}/\epsilon l_{B}$ in Figure 6). This is understandable since removing one electron creates three Laughlin quasiholes, but four MR quasiholes, and the latter leads to many more configurations with similar energies.
We emphasize that while detailed peak positions and splittings in the LDOS depend on the microscopic form of the interaction, certain qualitative features are universal. For example, in the Laughlin case, the emergence of two dominant peak clusters from the 3-quasihole spectrum is present across different interaction models and FWHM values, where the lower cluster contributes to the quasihole excitations. In the Moore–Read case, the LDOS shows a clear distinction between even and odd electron number sectors, a direct consequence of the underlying non-Abelian fusion rules. These features are expected to survive under reasonable experimental conditions and provide a possible path toward distinguishing topological signatures from non-universal effects in STM experiments.
@@ V Conclusion
In this paper, we investigated the effective interactions of anyons in Laughlin and Moore-Read FQH phases derived from diverse electron-electron interactions, both model and realistic ones. The effective interaction within anyons is an interpolation between fermions’ repulsive and bosons’ attractive force, exhibiting various and unusual behaviors depending on the detailed forms of the electron-electron interaction in the 2DEG. The recently developed STM experiments in FQH systems could be a new platform to explore these properties of anyons. As we show, many bound anyon states with different internal structures will be involved in the low-lying energy parts of the tunneling spectrum. We point out that effective anyon interactions are important factors in understanding this part of the results in STM tunneling experiments. Since anyons can be considered the elementary particles for the corresponding FQH phases and are potentially useful for the storage and processing of quantum information, the interplay between their effective interactions and the correlated topological phases should be considered as a significant feature when investigating their properties both theoretically and for future applications.
Note.
While this manuscript was under review, a related preprint appeared  [77] on the formation of anyon clusters (anyon “molecule”), readers may find it complementary.
@@ Acknowledgements.
Some of the numerics in this work are performed using the DiagHam package for which we would like to thank all the authors. This work is supported by the National Research Foundation, Singapore, under the NRF Fellowship Award (NRF-NRFF12-2020-005), Singapore Ministry of Education (MOE) Academic Research Fund Tier 3 Grant (No. MOE-MOET32023-0003) “Quantum Geometric Advantage”, and Singapore Ministry of Education (MOE) Academic Research Fund Tier 2 Grant (No. MOE-T2EP50124-0017).